OpenStudy (itsbribro):

help please given f(x)= -x+15, what is f(-3)? A. -45 B. 12 C. 18 D. 45

OpenStudy (acxbox22):

plug in -3 in the place of x and solve f(-3)=-(-3)+15 f(-3)=3+15 f(-3)=??

OpenStudy (itsbribro):

18

OpenStudy (acxbox22):

yes

OpenStudy (itsbribro):

thank you @acxbox22

OpenStudy (acxbox22):

np